Question Two

a) Suppose that you have an extra K111.3 today that you wish to invest for one year. If you can earn 11.1% per year on your investment, how much will you have in one year?
                                                                                                     (10 marks)
1.) Compared to investing in a single security, how does diversification help the investors?                                                                                   (3 marks)
2) In what situation is Portfolio diversification least likely to protect against losses:
3) Explain the features of a defined contribution pension plan: (5 marks)
• Employer-sponsored retirement plans
• As the names imply, a defined benefit pension plan provides a specified payment amount in retirement
• Employers guarantee a specific retirement benefit amount for each participant of a defined benefit plan which can be based on the employee's salary, years of service, or a number of other factors.
• Employees have little control over the funds until they are received in retirement.
• The employer bears the investment risk of ensuring the defined benefit amount is able to be paid to the retired employee
• Due to this risk, defined benefit plans require complex actuarial projections and insurance for guarantees, making the costs of administration very high
4) Explain why an insurance company has Low risk tolerance and high liquidity requirements’                                                                              (3 marks)
5) Explain why an endowment as long time horizon and low liquidity (4 marks)
Answer #1

Question Two

a)

fix =FV(B2,B3,0,-B1) В4 A 111.3 1 Current Amount 2 Rate 0.111 Period in years 1 3 4 Amount after 1 year =FV(B2,B3,0,-B1)

fox X В4 =FV( A В 1 Current Amount 111.3 2 Rate 11.10% 3 Period in years 4 Amount after 1 year 123.65 L

After 1 year, amount available will be K 123.65.

Note: Only 1 separate question per post is answered. Please ask the remaining questions in a separate post.